The Golden Valley baseball team played 28 games. They won three times the number of Games that they lost.How many games did they win? How many games did they lose?
Answer by TimothyLamb(4379)

W + L = 28
W = 3L
---
W + L = 28
3L + L = 28
4L = 28
---
L = 7
W = 21
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php